OpenStudy (anonymous):

Two marbles are drawn without replacement from a box with 3 white, 2 green, and 1 blue marble. Find the probability of drawing one white marble and one blue marble.

OpenStudy (anonymous):

your help is greatly appreciated

OpenStudy (kc_kennylau):

First you draw one white marble Then you'll be left with 2 white 2 green 1 blue Then you draw one blue marble

OpenStudy (anonymous):

I think 4/15 is the answer. Can you confirm?

OpenStudy (anonymous):

First, I'm assuming that the order in which the marbles are drawn is not important. Thus, the probability of drawing a white marble and a blue marble is: \[ = (\frac{ 3 }{ 6 })(\frac{ 1 }{ 5 }) + (\frac{ 1 }{ 6 })(\frac{ 3 }{ 5 }) = \frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

I Think the answer is \[\frac{ 1 }{ 10 }\] because the desired event is one white and one blue without replacement and the total number of balls is 6 then: \[\frac{ 3 }{ 6 } \frac{ 1 }{ 5 } = \frac{ 3 }{ 30 } = \frac{ 1 }{ 10 } \]

OpenStudy (anonymous):

Welcome to OpenStudy @josie2013 ! Here's a link to help you around OpenStudy! http://openstudy.com/study#/updates/52f5a913e4b0512c88e5ef0f